OpenStudy (anonymous):

A certain type of chocolate comes in a package that's the shape of a triangular prism. How much chocolate is contained in this package? triangular prism having a triangle height of 4 centimeters, a base of 3 centimeters, and a prism height of fourteen centimeters Answer one hundred thirty-eight cubic centimeters eighty-four cubic centimeters five hundred four cubic centimeters one hundred sixty-eight cubic centimeters

OpenStudy (inkyvoyd):

(4*3/2)*14=6*14=84

OpenStudy (anonymous):

thank you